Chlorine gas occupies a volume of 25 mL at 300K. What volume will it occupy at 600K? 



1
Expert's answer
2021-03-08T05:55:33-0500

"V_1 = 25 \;mL \\

T_1 = 300 \;K \\

T_2 = 600 \;K \\

P = constant \\

\frac{V_1}{T_1}= \frac{V_2}{T_2} \\

V_2 = \frac{V_1 \times T_2}{T_1} \\

= \frac{25 \times 600}{300} \\

= 50 \;mL"

Answer: 50 mL